One number is 12 more than another. The sum of the smaller number and twice the larger is 39. Find the larger number.

Answer by stanbon(75887)

One number is 12 more than another. The sum of the smaller number and twice the larger is 39. Find the larger number.
-----------
One number is "x" ; the other number is "x+12",
EQUATION:
x + 2(x+12) = 39
3x+24 = 39
3x = 15
x = 5 (one of the numbers)
x+12= 17 (the other number)
===================
Cheers,
Stan H.